40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3295     Entry time: Mon Jul 26 20:30:35 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: 2 Tip Tilts suspended and balanced 

[Koji, Jenne]

We were on Team Cleanroom, while Kiwamu and Alberto were on Team Chamber.  Team Cleanroom suspended and balanced 2 Tip Tilts this afternoon.

One of the TTs that was suspended today is the one which was broken on Friday (see elog 3278).  We resuspended it using the regular 0.0036" diameter wire (91um).  We balanced it using the HeNe oplev, and then set it aside.  This TT has serial number 2.

We noticed that, like the previous 2 TT suspensions (this one before it was broken, and the one actually installed in the BS chamber on Friday, which is #3), there seems to be a little bit of hysteresis in the pointing.  The difference comes if we poke the top of the mirror holder and observe the place the reflected beam spot comes to rest at, and if we poke the bottom of the mirror holder.  The beam spot stays a little higher when we poke the top vs. when we poke the bottom. 

To combat this, we tried suspending our second TT of the day (the one that Kyung Ha and I had half finished) using thinner wire for the mirror holder.  We used the 0.0017" diameter wire (43um) that is used for the SOSes.  Unfortunately, it still seems like there is a similar hysteresis.  The thin-wire TT has serial number 4.

While working on TT4, we recalled that we have to include rubber dampers for the vertical blade springs.  Oooops!  We used some of the leftover #4-40 screws with viton tips that Zach and Mott had made for Earthquake stops to damp the vertical resonance of the blades.  We measured the Q factor by flicking the blades up or down.  We changed the oplev setup to be a shadow sensor setup, and watched the ringdown of the vertical mode on the 'scope.  We counted #cycles/time = frequency, and the t(1/2) time for the exponential ringdown to calculate the Q.  For the shadow sensor, we positioned the QPD in line with the initial HeNe beam, and placed the edge of the mirror holder clamp partially in the beam, so the beam was partly occluded.  When the mirror shook up and down, more or less of the beam was blocked, and we could see this power fluctuation on the 'scope.

Using the formula Q = pi  f0 T1/2 / ln(2) = 4.53 f0 T1/2, where T1/2 is the the time it takes for the amplitude to decay by half, we measured a Q of 31 for the vertical mode with no damping, and a Q of 14 with damping.  Koji confirmed the calculation and put it into wiki.

We need to go through the other TTs that have been assembled and give them their rubber dampers.

 

ELOG V3.1.3-